# Self-intersection Number of Negative Curves on Fermat Surfaces

### v1

@@ Abstract.
We give an explicit formula for the self-intersection number of negative curves on Fermat surfaces. The formula gives us hints to either prove or disprove the Bounded Negativity Conjecture for the Fermat surfaces.
@@ 1. Introduction
The Bounded Negativity Conjecture holds on the Fermat surface $X_{d}$ if one of the following conditions is satisfied:
@@
On the other hand, the formula in Theorem 1.1 also gives us hint to construct counterexamples for the Bounded Negativity Conjecture on Fermat surfaces; see Section 5.
@@
In particular, for $d\leq 4$, the bounded negativity of $X_{d}$ follows from this argument, since the canonical divisor of $X_{d}$ is $K_{X_{d}}=(d-4)H$ where $H$ is the hyperplane section. However, for $d\geq 5$, the canonical divisor $K_{X_{d}}$ is ample and $X_{d}$ is of general type. To the author’s knowledge, there are currently no effective methods available to deal with the Bounded Negativity of such surfaces.
Another natural approach to establishing the Bounded Negativity for a surface is to show that there are only finitely many negative curves. However, it is proved in [3, Thm.4.1] that for any $m>0$, there exists a smooth projective surface which admits infinitely many smooth irreducible curves of self-intersection number $-m$. For the Fermat surface $X_{d}$, we do not know whether the number of negative curves on $X_{d}$ is finite or not.
@@
Other strategies to prove the Bounded Negativity Conjecture include investigating the non-isomorphic endomorphisms or establishing dominance by other surfaces with bounded negativity [3, Prop.2.1], exploring the birational invariance of bounded negativity and the $H$-constant [2], using the equivalence between bounded negativity and the boundedness of the denominators in the Zariski decomposition [4], discussing the weighted Bounded Negativity Conjecture [11, 8], and studying the connections between the Bounded Negativity Conjecture and other unsolved problems, such as SHGH conjecture or the bounded cohomology conjecture [1, 7]. However, despite providing invaluable insights in understanding the Bounded Negativity Conjecture, these methods either deal with only very special cases, e.g., rational surfaces or K3 surfaces, or just offer general frameworks that reduce the conjecture to other open questions. They rely on general formulas, such as the adjunction formula and Riemann-Roch formula, to establish bounds for the intersection number $C^{2}$ of a reduced and irreducible curve $C$, without using specific information about $C$, such as its equation.
@@
For any $d\geq 1$, there exists a constant $c_{d}$ that depends only on $d$ such that all smooth surfaces of degree $d$ in $\mathbb{P}^{3}$ has negativity bound $c_{d}$.
@@
Hence, to establish the Bounded Negativity Conjecture for surfaces in $\mathbb{P}^{3}$, it suffices to prove its validity for a specific surface. Our Theorem 1.1 is proved for this purpose.
The formulas in Theorem 1.1 show how the Bounded Negativity Conjecture is connected with the singularity theory of curves. The key term in the expression of the self-intersection number is the sum $\sum\limits_{p\in D\cap E}\mu_{p}(E)$. It is notable that this sum is taking over the intersection points of $D$ and $E$, not all the singular points of $E$.
To conclude the introduction, we give an outline of the proof of Theorem 1.1. Given a negative curve $C$ on $X_{d}$. First, we represent the Fermat surface $X_{d}$ as a branched Galois cover of $\mathbb{P}^{2}$:
$\rho:\ X_{d}\to\mathbb{P}^{2}\quad[x,y,z,w]\mapsto[x,y,z]$
Let $E=\rho(C)$ and $\mathscr{C}=\rho^{*}E$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$; denote by $g$ its defining equation. It is clear that $g$ can be seen as a homogeneous polynomial on $\mathbb{P}^{3}$ which vanishes on $C$. Second, we prove that $\mathscr{C}$ is a reduced curve on $X_{d}$; see Proposition 4.4. Moreover, $\mathscr{C}$ is of the following form
@@
with $C_{i}^{2}=C$ for all $i=1,\cdots,k$. This result holds because $C_{i}$ lie in the orbit of $C$ under the Galois group action for $\rho$. It follows that
@@
Hence, to compute $C^{2}$, we need to calculate the sum on the right hand side. Third, the intersection number can be calculated in a local way at any intersection point. Locally, the curve $\mathscr{C}$ can be represented as the curve in $(\mathbb{C}^{3},0)$ defined by
@@
We will first solve $v$ as Puiseux series of $u$ from the first equation, and then use $l\left(u,v(u)\right)+\omega^{d}=0$ to investigate the intersection multiplicities between different irreducible components of $\mathscr{C}$. This is the most technical part in this paper and is done in Proposition 4.4 and Section 6.
The author would like to thank the organizers of International Symposium on Singularities and Applications, Dec. 9-13, 2024 in Sanya, China. The stimulating atmosphere of the conference inspired him to conceive the idea of performing computation using Puiseux series. The author was partially supported by the grant YSBR-001 and is supported by NSFC No. 12301100.
@@ 2. Curves on surfaces of Sebastiani-Thom type
Assume that $X\subset\mathbb{P}^{3}$ be a smooth surface defined by the following equation:
@@
where $f$ is a binary form of degree $d$. Such a surface is said to be of Sebastiani-Thom type. The Fermat surface is a typical example. Let $C$ be a reduced and irreducible curve on $X$.
@@
    where $C_{1}=C,C_{2},\cdots,C_{k}$ are distinct reduced and irreducible curves and $m\geq 1$ is an integer.
@@
    The self-intersection numbers $C_{i}^{2},i=1,\cdots,k$ are all equal to $C^{2}$.
@@
Let $g(x,y,z)$ be an irreducible defining equation for $E$. Then the inverse image $\rho^{-1}(E)$ is exactly the curve cut out by $g$ on $X$. We may see $g$ as a homogeneous polynomial on $X$ and write the divisor $(g)$ cut out by $g$ on $X$ as
@@
where $C_{1},\cdots,C_{k}$ are reduced and irreducible curves on $X$ and $m_{1},\cdots,m_{k}\geq 1$.
Since the action $\mathbb{Z}_{d}$ for the morphism $\rho$ is a Galois action, we see that $\rho^{-1}(E)$ is exactly the union of the curves in the $\mathbb{Z}_{d}$-orbit of $C$. It follows that $C_{i},i>1$ are all isomorphism to $C$ by an automorphism of $X$. Hence $C_{i}^{2}=C^{2}$ for all $i=1,\cdots,k$; and (i) follows.
Moreover, we may see $g$ as a $\mathbb{Z}_{d}$-invariant polynomial. For each pair $i\neq j$, there exists an element $\alpha$ in the Galois group $\mathbb{Z}_{d}$ such that $\alpha\cdot C_{i}=C_{j}$, and then
@@ 3. Strategy of Proof
Let $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$ is the Fermat surface in $\mathbb{P}^{3}$ and $C\subset X_{d}$ is a reduced and irreducible curve. We shall investigate the lower bound of the self-intersection number $C^{2}$.
@@
Then $C_{1},\cdots,C_{k}$ are the irreducible components of $\mathscr{C}$.
@@ 3.1. Strategy
where $C_{i}\cdot C_{j}$ is the intersection number on $X_{d}$.
To investigate the Bounded Negativity Conjecture on $X_{d}$, we need to obtain a lower bound for $C^{2}$. From the formula above, we are led to give an upper bound for the sum
@@
The result will be a function in $e$ where $d$ is viewed as a constant, and so is the difference
@@
The main part of this paper is to calculate the intersection numbers by using local computation.
For an arbitrary intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of some $C_{i}$ and $C_{j}$ with $i\neq j$, we will give an explicit formula for the sum
@@
where $I_{P}(C_{i},C_{j})$ is the intersection multiplicity (that will be explored in detail in Section 6).
Since we consider all curves $C$ on $X_{d}$, we see $g$ as an abstract polynomial. From $P\in X_{d}$, we have that $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}+w_{0}^{d}=0$, hence not all $x_{0},y_{0},z_{0}$ are zero. By the symmetry of the coordinates $x,y,z$ in the equation $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$, we may without loss of generality assume that $x_{0}\neq 0$. To decide where the intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ may situate on $X_{d}$, we propose the following Lemma which will be proved later.
@@
where $w$ is the fourth homogeneous coordinate for $\mathbb{P}^{3}$. Moreover, $Q\in C_{i}$ for all $i=1,\cdots,k$, i.e., $Q$ lies on every irreducible component of $\mathscr{C}$.
@@
Then the curve $\mathscr{C}:g=0$ on $X_{d}$ is defined in $\mathbb{C}^{3}=\{x\neq 0\}\subset\mathbb{P}^{3}$ by
@@
$\frac{\partial\left(l(u,v)+\omega^{d}\right)}{\partial v}(0,0,0)=d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}\neq 0.$
@@
Our method is, however, first solving $v$ from the equation $h(u,v)=0$ by using the Puiseux series method (that will be recalled in Section 6), say $v=s(u)$. Then using the equation $l(u,s(u))+\omega^{d}=0$ to discuss the properties of the curve $\mathscr{C}$ and calculate the intersection multiplicities. To this end, we need that $h$ is $v$-general (the related notions will be discussed in Section 6).
At this point, we specially consider the exceptional case where $h$ is not $v$-general, or equivalently, $h(0,v)\equiv 0$; the $v$-general case will be investigated in subsequent sections.
Recall that $h$ is transformed from an irreducible polynomial $g$, thus $h$ is an irreducible polynomial in $u,v$. If $h$ is not $v$-general, then $u$ is a factor of $h$ and so $h$ is a constant multiple of $u$ by irreducibility. Back to $g$, we see that $g$ is a linear function $y-ax$ (up to multiplicative factor). Therefore, the curve $\mathscr{C}$ is defined as
@@
Recall that we used the degree $d$ branched Galois cover $\rho:X_{d}\to\mathbb{P}^{2}$ defined by $[x,y,z,w]\mapsto[x,y,z]$. The ramification locus of $\rho$ is the closed subset of $X_{d}$ defined by the equation $w=0$.
If $Q$ is a point on at least two components of $\mathscr{C}$, then #$\#\left(\rho^{-1}(\rho(Q))\right)<d$, where #$\#A$ means the number of points of a set $A$. It follows that $Q$ lies on the ramification locus of $\rho$. Thus $w(Q)=0$ as desired.
@@ 4.1. Position of intersection points
If $q=[x,y,z]$ lies on both $D$ and $E$, then $Q=[x,y,z,0]$ lies on all of $C_{i},i\leq k$.
@@
By definition, $\rho^{-1}(q)=Q$ and $Q\in\mathscr{C}$. Since each $C_{i}$ is mapped onto $E$, the intersection $\rho^{-1}(q)\cap C_{i}$ is nonempty. Hence $Q\in C_{i}$ as desired.
@@ 4.3. Intersection multiplicities
To give a lower bound for $C^{2}$, we need to estimate the sum on the right hand side. Moreover, by definition, we have
$\sum_{i\neq j}C_{i}\cdot C_{j}=\sum_{\stackrel{{\scriptstyle i\neq j}}{{Q\in C_{i}\cap C_{j}}}}I_{Q}(C_{i},C_{j}).$
Hence, it suffices to do computation for the sum
@@
where $P=[x_{0},y_{0},z_{0},0]$ is chosen above. Our main technical results are as follows.
@@
    The following equation holds:
@@
To investigate the Bounded Negativity Conjecture for Fermat surfaces, we only need to consider negative curves $C$. Then the curve $\mathscr{C}$ necessarily have $k\geq 2$ irreducible components since otherwise $C^{2}$ would be nonnegative. Moreover, $k$ divides $d$ since the Galois action for $\rho$ is a cyclic action of order $d$. To ensure that the polynomial $h$ is $v$-general, we only need to assume $C^{2}\neq 2-d$. Therefore, combining all of our discussions above, we get the following theorem.
@@
$g(E)=\frac{(e-1)(e-2)}{2}-\sum_{p\in E}\left(\mu_{p}(E)+r_{p}-1\right)$
by [12, Cor.7.1.3], where $r_{p}$ is the number of local branches of $E$ at $p$. If follows that
@@
for some $\alpha>0$. The result follows by noting that

### v2

@@ Abstract.
We give an explicit formula for the self-intersection number of negative curves on Fermat surfaces. The formula offers us hints to either prove or disprove the Bounded Negativity Conjecture for the Fermat surfaces.
@@ 1. Introduction
The Bounded Negativity Conjecture holds for the Fermat surface $X_{d}$ if one of the following conditions is satisfied:
@@
for some constant $\lambda>0$ which may depend on $d$, where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$; or
@@
On the other hand, the formula in Theorem 1.1 also gives us hints to construct counterexamples for the Bounded Negativity Conjecture on Fermat surfaces; see Section 5.
@@
In particular, for $d\leq 4$, the bounded negativity of $X_{d}$ follows from this argument, since the canonical divisor of $X_{d}$ is $K_{X_{d}}=(d-4)H$ where $H$ is the hyperplane section. However, for $d\geq 5$, the canonical divisor $K_{X_{d}}$ is ample and $X_{d}$ is of general type. To the author’s knowledge, there are currently no effective methods available to deal with the Bounded Negativity of such surfaces, except for the weak Bounded Negativity Conjecture which will be quoted below.
Another natural approach to establishing the Bounded Negativity for a surface is to show that there are only finitely many negative curves on the surface. However, it is proved in [3, Thm.4.1] that for any $m>0$, there exists a smooth projective surface which admits infinitely many smooth irreducible curves of self-intersection number $-m$. For the Fermat surface $X_{d}$, we do not know whether the number of negative curves on $X_{d}$ is finite or not.
@@
Other strategies to prove the Bounded Negativity Conjecture include investigating the non-isomorphic surjective endomorphisms or establishing dominance by other surfaces that have bounded negativity [3, Prop.2.1], exploring the birational invariance of bounded negativity and the $H$-constant [2], using the equivalence between bounded negativity and the boundedness of the denominators in the Zariski decomposition [4], discussing the weighted Bounded Negativity Conjecture [11, 8], and studying the connections between the Bounded Negativity Conjecture and other unsolved problems, such as SHGH conjecture or the bounded cohomology conjecture [1, 7]. However, despite providing invaluable insights in understanding the Bounded Negativity Conjecture, these methods either deal with only very special cases, e.g., rational surfaces or K3 surfaces, or just offer general frameworks that reduce the conjecture to other open questions. They rely on general formulas, such as the adjunction formula and Riemann-Roch formula, to establish bounds for the intersection number $C^{2}$ of a reduced and irreducible curve $C$, without using specific information about $C$, such as its equation.
@@
For any $d\geq 1$, there exists a constant $c_{d}$ that depends only on $d$ such that all smooth surfaces of degree $d$ in $\mathbb{P}^{3}$ have negativity bound $c_{d}$.
@@
Assuming the conjecture above, to establish the Bounded Negativity Conjecture for surfaces in $\mathbb{P}^{3}$, it suffices to prove its validity for a specific surface. Our Theorem 1.1 is proved for this purpose.
The formulas in Theorem 1.1 show how the Bounded Negativity Conjecture is connected with the singularity theory of curves. The key term in the expression of the self-intersection number $C^{2}$ is the sum $\sum\limits_{p\in D\cap E}\mu_{p}(E)$. It is notable that this sum is taken over the intersection points of $D$ and $E$, not all the singular points of $E$.
To conclude the introduction, we give an outline of the proof of Theorem 1.1. Given a negative curve $C$ on $X_{d}$, we represent the Fermat surface $X_{d}$ as a branched Galois cover of $\mathbb{P}^{2}$:
$\rho:\ X_{d}\to\mathbb{P}^{2},\quad[x,y,z,w]\mapsto[x,y,z]$
Let $E=\rho(C)$ and $\mathscr{C}=\rho^{*}E$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$; denote by $g$ its defining equation. It is clear that $g$ can be seen as a section of the line bundle $\mathscr{O}_{X_{d}}\left((\deg g)H\right)$ which vanishes on $C$. Then we prove that $\mathscr{C}$ is a reduced curve on $X_{d}$; see Proposition 4.4. Moreover, $\mathscr{C}$ is of the following form
@@
with $C_{i}^{2}=C$ for all $i=1,\ldots,k$. This result holds because $C_{i}$’s lie in the orbit of $C$ under the Galois group action for $\rho$. It follows that
@@
Hence, to compute $C^{2}$, we need to calculate the sum on the right hand side. Finally, the intersection numbers can be calculated in a local way at any intersection point. Locally, the curve $\mathscr{C}$ can be represented as the curve in $(\mathbb{C}^{3},0)$ defined by
@@
We will first solve $v$ as Puiseux series of $u$ from the first equation, and then use $l\left(u,v(u)\right)+\omega^{d}=0$ to investigate the intersection multiplicities between different irreducible components of $\mathscr{C}$. This is the most technical part in this paper and will be completed in Proposition 4.4 and Section 6.
The author would like to thank the organizers of International Symposium on Singularities and Applications, Dec. 9–13, 2024 in Sanya, China. The stimulating atmosphere of the conference inspired him to conceive the idea of performing computation using Puiseux series. The author was partially supported by the grant YSBR-001 and is currently supported by NSFC No. 12301100.
@@ 2. Curves on surfaces of Sebastiani-Thom type
Assume that $X\subset\mathbb{P}^{3}$ is a smooth surface defined by the following equation:
@@
where $f$ is a form of degree $d$. Such a surface is said to be of Sebastiani-Thom type. The Fermat surface is a typical example. Let $C$ be a reduced and irreducible curve on $X$.
@@
    where $C_{1}=C,C_{2},\ldots,C_{k}$ are distinct reduced and irreducible curves and $m\geq 1$ is an integer.
@@
    The self-intersection numbers $C_{i}^{2},i=1,\ldots,k$ are all equal to $C^{2}$.
@@
Let $g(x,y,z)$ be an irreducible defining equation for $E$. Then the inverse image $\rho^{-1}(E)$ is exactly the curve defined by $g=0$ on $X$ (here we see $g$ as a homogeneous polynomial in $x,y,z,w$ restricted to $X$). Write the divisor $(g)$ cut out by $g$ on $X$ as
@@
where $C_{1},\ldots,C_{k}$ are reduced and irreducible curves on $X$ and $m_{1},\ldots,m_{k}\geq 1$.
Since the action $\mathbb{Z}_{d}$ for the morphism $\rho$ is a Galois action, we see that $\rho^{-1}(E)$ is exactly the union of the curves in the $\mathbb{Z}_{d}$-orbit of $C$. It follows that $C_{i},i>1$ are all isomorphic to $C$ by automorphisms of $X$. Hence $C_{i}^{2}=C^{2}$ for all $i=1,\ldots,k$; and (i) follows.
Moreover, it is obvious that $g$ is a $\mathbb{Z}_{d}$-invariant polynomial in $(x,y,z,w)$. For each pair $i\neq j$, there exists an element $\alpha$ in the Galois group $\mathbb{Z}_{d}$ such that $\alpha\cdot C_{i}=C_{j}$, and then
@@ 3. Strategy of Proof
Let $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$ be the Fermat surface in $\mathbb{P}^{3}$ and $C\subset X_{d}$ a reduced and irreducible curve. We shall investigate the lower bound of the self-intersection number $C^{2}$.
@@
Then $C_{1},\ldots,C_{k}$ are the irreducible components of $\mathscr{C}$.
@@ 3.1. Strategy
where $C_{i}\cdot C_{j}$ denotes the intersection number on $X_{d}$.
To study the Bounded Negativity Conjecture on $X_{d}$, we need to obtain a lower bound for $C^{2}$. From the formula above, we are led to give an upper bound for the sum
@@
The result will be a function of $e$ where $d$ is viewed as a constant, and so is the difference
@@
The main part of this paper is to calculate the intersection numbers using local computations. The results are summarized in Proposition 4.4, and the lengthy proof will be presented in Section 6.
Fix an arbitrary intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of some $C_{i}$ and $C_{j}$ with $i\neq j$, we will give an explicit formula for the sum
@@
where $I_{P}(C_{i},C_{j})$ denotes the intersection multiplicity (that will be explored in detail in Section 6).
Since we consider all curves $C$ on $X_{d}$, we should view $g$ as an abstract polynomial. From $P\in X_{d}$, we have that $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}+w_{0}^{d}=0$, hence not all $x_{0},y_{0},z_{0}$ are zero. By the symmetry of the coordinates $x,y,z$ in the equation $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$, we may without loss of generality assume that $x_{0}\neq 0$. To decide where the intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ may situate on $X_{d}$, we propose the following Lemma which will be proved at the end of this section.
@@
where $w$ is the fourth homogeneous coordinate for $\mathbb{P}^{3}$. Moreover, $Q\in C_{i}$ for all $i=1,\ldots,k$, i.e., $Q$ lies on every irreducible component of $\mathscr{C}$.
@@
Then the curve $\mathscr{C}:g=0$ on $X_{d}$ is defined in the affine piece $\mathbb{C}^{3}=\{x\neq 0\}\subset\mathbb{P}^{3}$ by
@@
$\frac{\partial\left(l(u,v)+\omega^{d}\right)}{\partial v}\Bigg{|}_{(0,0,0)}=d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}\neq 0.$
@@
Our method is, however, first to solve $v$ from the equation $h(u,v)=0$ by using the Puiseux series method (which will be recalled in Section 6), say $v=s(u)$. Then, using the equation $l(u,s(u))+\omega^{d}=0$, we discuss the properties of the curve $\mathscr{C}$ and calculate the intersection multiplicities. To this end, we need that $h$ is $v$-general (the related notions will be discussed in Section 6).
At this point, we specially consider the exceptional case where $h$ is not $v$-general, or equivalently, $h(0,v)\equiv 0$; the $v$-general case will be investigated in Section 6.
Recall that $h$ is transformed from an irreducible polynomial $g$, thus $h$ is an irreducible polynomial in $u,v$. If $h$ is not $v$-general, then $u$ is a factor of $h$; therefore, by irreducibility, $h$ is a constant multiple of $u$. Back to $g$, we see that $g$ is a linear function $y-ax$ (up to multiplicative factor). Therefore, the curve $\mathscr{C}$ is defined as
@@
Recall that we have the degree $d$ branched Galois cover $\rho:X_{d}\to\mathbb{P}^{2}$ defined by $[x,y,z,w]\mapsto[x,y,z]$. The ramification locus of $\rho$ is the closed subset of $X_{d}$ defined by the equation $w=0$.
If $Q$ is a point on at least two components of $\mathscr{C}$, then #$\#\left(\rho^{-1}(\rho(Q))\right)<d$, where #$\#A$ means the number of points in a set $A$. It follows that $Q$ lies on the ramification locus of $\rho$. Thus, $w(Q)=0$ as desired.
@@ 4.1. Position of intersection points
If $q=[x,y,z]$ lies on both $D$ and $E$, then $Q=[x,y,z,0]$ lies on all of $C_{i}$ for $i\leq k$.
@@
By definition, $\rho^{-1}(q)=\{Q\}$ and $Q\in\mathscr{C}$. Since each $C_{i}$ is mapped onto $E$, the intersection $\rho^{-1}(q)\cap C_{i}$ is nonempty. Hence $Q\in C_{i}$ as desired.
@@ 4.3. Intersection multiplicities
To give a lower bound for $C^{2}$, we need to estimate the sum on the right-hand side. Moreover, by definition, we have
$\sum_{i\neq j}C_{i}\cdot C_{j}=\sum_{\begin{subarray}{c}i\neq j\\ Q\in C_{i}\cap C_{j}\end{subarray}}I_{Q}(C_{i},C_{j}).$
Hence, it suffices to compute the sum
@@
where $P=[x_{0},y_{0},z_{0},0]$ is chosen above. Our main technical results are summarized as follows.
@@
    The following equality holds:
@@
Now we use the above results to study the Bounded Negativity Conjecture for Fermat surfaces. It is clear that we only need to consider negative curves $C$. Then the curve $\mathscr{C}$ necessarily has $k\geq 2$ irreducible components since otherwise $C^{2}$ would be nonnegative. Moreover, the number $k$ divides $d$ since the Galois action for $\rho$ is a cyclic action of order $d$. To ensure that the polynomial $h$ is $v$-general, we only need to assume $C^{2}\neq 2-d$; see the discussions in Section 3. Therefore, combining all of our discussions above, we get the following theorem.
@@
where again $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is as defined above.
@@
$g(E)=\frac{(e-1)(e-2)}{2}-\sum_{p\in E}\frac{1}{2}\left(\mu_{p}(E)+r_{p}-1\right)$
by [12, Cor.7.1.3], where $r_{p}$ is the number of local branches of $E$ at $p$. It follows that
@@
for some $\alpha>0$. The desired result follows by noting that